# Upgrade of Super-Kamiokande with Gadolinium

### v1

@@ 1 INTRODUCTION
While many important measurements have been made with these WC detectors, a major drawback has been their very inefficient response to the presence of thermal neutrons. Now, an affordable and effective technique has been developed and put into practice at Super-Kamiokande to overcome this situation via the addition to water of an optically transparent, low background, low reactivity solute with a large neutron cross section and energetic gamma daughters: gadolinium sulfate octahydrate ($\rm Gd_{2}(\rm SO_{4})_{3}\cdot\rm 8H_{2}O$). This is illustrated in Figure 1.
@@
The idea of loading the famous Super-Kamiokande neutrino and nucleon detector with tons of a water soluble gadolinium compound, a concept which would initially be called GADZOOKS! [Gadolinium Antineutrino Detector Zealously Outperforming Old Kamiokande, Super![1] and then later officially put into practice as SK-Gd, was first introduced at the November 2002 SK Collaboration meeting. Almost immediately following this first public airing of the idea, studies began in Japan and the US to better understand the possible physics benefits, as well as the challenges, of this potential upgrade.
Since then, a lot has happened. Studies progressed. Vital pieces of equipment were invented and continually improved. More people (and nations) got involved. After investigating the use of gadolinium chloride and gadolinium nitrate as our Gd-containing solute, we settled on gadolinium sulfate as the best possible compound for use in Super-K due to its combination of high solubility and high transparency coupled with low reactivity with detector components. The year 2009 saw the funding and construction of EGADS [Evaluating Gadolinium’s Action on Detector Systems], a dedicated, large-scale test facility for gadolinium R&D [2]. This new underground laboratory was built in the Mozumi mine explicitly to demonstrate that the loading of Super-Kamiokande would be safe and effective (see Section 3).
@@ 2.1 Neutron Capture in Water
In pure water, the emitted neutrons thermalize and are captured by protons, with the newly formed deuterons each emitting a single 2.2 MeV gamma ray which can then be measured as delayed signals. Since the fourth phase of SK [SK-IV] started in 2008, new electronics equipment has been introduced [7], making it possible to measure delayed signals. However, in order to achieve the background reduction required for the DSNB analysis mentioned below, the neutron detection efficiency had to be set relatively low (to about 20%). But the situation changes with the introduction of gadolinium. When a neutron is captured by gadolinium, it emits a cascade of gamma rays with a total energy of about 8 MeV. This is higher than the energy threshold in SK, and thus the detection efficiency improves drastically even under the same trigger conditions as in the pure-water case. Furthermore, gadolinium has an exceptionally high neutron capture cross section compared to all other elements; it will capture about 75% of thermal neutrons with a modest concentration of only 0.03% Gd by mass dissolved in the water.
@@ 2.2 Diffuse Supernova Neutrino Background
Core-collapse supernovae are powerful explosions that occur in the final stages of the evolution of massive stars, during which roughly 99% of their enormous release of gravitational energy is carried away by neutrinos. Since the beginning of the universe, there have been o($10^{18}$) core-collapse supernovae, and the neutrinos released in these explosions fill the current universe as background radiation. This is called the diffuse supernova neutrino background [DSNB]. DSNB observations are expected to provide valuable insights into astrophysics, such as past star formation rates, the frequency of supernova explosions, and the formation of black holes which cannot be seen by optical observations [8, 9, 10, 11]. Since the 1980s, DSNB searches have been conducted using various detectors; however, before the start of SK-Gd no nonzero DSNB flux had been observed. Super-Kamiokande currently has the most stringent constraints on this flux, but the expected event rate of DSNB interactions even in the huge SK detector is quite low - at most a few per year - and they are buried under background events, making observation extremely difficult.
@@ 2.3 Galactic Supernova Neutrinos
As mentioned above, many supernovae have occurred to date, but supernova neutrinos have only been observed once. In 1987, neutrinos from a supernova explosion that occurred in the Large Magellanic Cloud, 50 kpc from Earth, were observed at Kamiokande-II, IMB, and Baksan, with 11, 8, and 5 neutrino events, respectively. This observation proved that the basic mechanism of gravitational core collapse that occurs in the final stages of star evolution is correct. Unfortunately, the rate of supernovae close enough to detect neutrinos is estimated at only a few per century [12], so the chance of one occurring within the next decade is low, around 10-30%. However, if such a supernova occurs and a large number of neutrino events are observed, it would be a great advance in our understanding of core-collapse supernovae. Therefore, neutrino observatories around the world are currently preparing for the next nearby supernova to occur.
If a supernova occurs at the center of our galaxy, Super-Kamiokande will detect thousands of neutrinos. Here, the dominant neutrino interaction is the IBD described above, but if a large number of neutrinos are detected, we can also expect detection through other interactions. These include, for example, elastic scattering [ES] : $\nu+e^{-}\rightarrow\nu+e^{-}$; charged current interaction on oxygen : $\nu_{e}+\rm{{}^{16}O}\rightarrow e^{-}+\rm{{}^{16}F}$, $\bar{\nu_{e}}+\rm{{}^{16}O}\rightarrow e^{+}+\rm{{}^{16}N}$; and neutral current interaction on oxygen : $\nu+\rm{{}^{16}O}\rightarrow\nu+\rm{{}^{16}O^{*}}\rightarrow\nu+\rm{{}^{15}O/^{15}N}+\rm{n/p}+\gamma$. In particular, the direction of electrons produced by ES is strongly correlated with the direction of the original neutrinos, and therefore ES events can be used by Super-Kamiokande to determine the location of the supernova in the sky. Here, it is extremely important to determine the direction to the supernova as quickly as possible. Neutrinos are made before the huge increase in brightness of the dying star, and thus these neutrinos arrive at Earth earlier – anywhere from minutes to hours, depending on the stellar environment and type of progenitor star – than the supernova’s electromagnetic signals. Consequently, if the sky location of an ongoing explosion obtained from neutrino observations can be transmitted quickly enough to the astronomical community, then it should be possible for them to observe the beginning of the shock breakout of a core-collapse supernova.
@@ 2.4 Pre-supernova Neutrinos
As mentioned in the previous section, many neutrinos are expected to be seen by SK-Gd when a core-collapse supernova occurs within our galactic neighborhood, but if the explosion is relatively close to Earth (<<1 kpc) neutrinos could also be observed just before core collapse. Stars during the final stage of their lives are supported by nuclear fusion of heavy nuclei such as silicon, and the cooling mechanism is mainly by neutrino emission via thermal and nuclear processes. These are called pre-supernova neutrinos. Although pre-supernova neutrinos have not yet been found, their observation is very important because this silicon burning phase is thought to last for several days before the core collapse; detecting these neutrinos opens up the possibility of creating a supernova alert. It also would serve to guard against the unfortunate case of SK being down for planned calibration or maintenance just when a nearby star explodes. However, because the flux of these pre-supernova neutrinos is low, the possibility of their measurement is limited to very large, very nearby stars such as Betelgeuse. In fact, Super-Kamiokande has achieved highly efficient neutron capture in SK-Gd and significantly reduced background events, making the detection of neutrinos before a supernova explosion a realistic goal. Since 2021, SK-Gd has been monitoring pre-supernova neutrinos. The sensitivity of the system strongly depends on the supernova model and mass ordering, but in the case of Betelgeuse, for example, it should be possible to issue an alarm between 2 and 12 hours before its core collapse at a significance level corresponding to a false alarm rate of no more than 1 per century. In addition, joint observations with the KamLAND experiment began in May 2023. By integrating the complementary characteristics of the two detectors, the combined alarm system improves its sensitivity to pre-supernova neutrinos. For example, in the most optimistic model, the warning time is 1.5 hours longer than with SK-Gd alone. The combined alert system also reduces the dead time for pre-supernova neutrino detection and facilitates continuous monitoring even if one of the detectors is temporarily offline. [^1] As is currently the case for KamLAND at the time of this writing. That detector is temporarily offline for a major upgrade that started in 2024 and is expected to be completed in 2027.
@@ 2.5.1 Reactor Neutrinos
One of the physics opportunities of SK-Gd is the observation of reactor neutrinos. Since the spectrum of reactor neutrinos is soft, the detection is overwhelmed by radioactive backgrounds and hopeless if we use only the prompt signals of IBD. However, if a delayed neutron signal can be detected, the background events can be drastically reduced, opening the possibility of observing reactor neutrinos. Nevertheless, it is necessary to realize a lower energy threshold for the prompt signal as compared to the DSNB search. This is a long-baseline reactor neutrino observation, famous for measuring neutrino oscillations for $\theta_{12}$ and $\Delta m^{2}_{21}$ in the KamLAND experiment. In SK-Gd, reactor neutrinos are expected to be observed on the order of 100 per year from the currently limited cohort of operating Japanese reactors.
@@ 2.5.2 Atmospheric Neutrinos
In GeV-order neutrino observations, improving the neutron detection efficiency using Gd is expected to improve the ability to distinguish between neutrinos and antineutrinos. This is because antineutrinos are more likely to produce neutrons in the primary interaction than neutrinos, and the number of neutrons detected after neutrino interactions is expected to reflect this tendency. However, the number of neutrons produced in the final state because of interactions within the nucleus increases with energy, and the neutron multiplicity is much larger than would be predicted simply by considering only primary interactions. Therefore, understanding hadron-nucleus interactions inside the nucleus and in the detector is crucial for this purpose. For atmospheric neutrinos, the ability to distinguish between neutrinos and antineutrinos improves the sensitivity to neutrino mass ordering by taking advantage of the different oscillation patterns of their neutrinos as they pass through the Earth’s core and mantle.
@@ 2.5.3 Accelerator Neutrinos
T2K is a long-baseline neutrino experiment in which accelerator neutrinos generated at J-PARC in Tokai, Ibaraki, Japan, are observed at SK, approximately 300 km away. This experiment is expected to also improve its measurement sensitivity by improving the ability to distinguish between neutrinos and antineutrinos, as mentioned in the previous section. Increased neutron detection efficiency is particularly important for antineutrino appearance analysis, as contamination of neutrino events is currently expected to be relatively large. Preliminary studies have shown that this contamination is reduced from 30% to 13% by gadolinium loading. In this study, we have confirmed that, although still very preliminary, the introduction of additional $\bar{\nu}$-enhanced and $\nu$-enhanced samples slightly improves the sensitivity in the analysis of CP violation in the lepton sector.
@@ 3 EGADS DEMONSTRATOR
The EGADS facility includes a 200 ton scale model of SK complete with 240 photomultiplier tubes (227 50-cm SK style tubes plus an additional 13 prototype tubes which were evaluated for use in the Hyper-Kamiokande project), and two different selective water filtration systems. There is also a 15-ton pre-mixing and pre-treatment tank where the $\rm Gd_{2}(\rm SO_{4})_{3}\cdot\rm 8H_{2}O$is dissolved and cleaned prior to injection into the main 200-ton tank, and a custom-built water transparency measuring system called UDEAL [Underground Device Evaluating Attenuation Length].
@@ 4 TANK REFURBISHMENT FOR GD-LOADING
Since the beginning of Super-K’s operation in 1996, there has been a water leak from the Super-K tank, ranging from 1 to 4 tons per day, depending on the water level in the surrounding rock. That amount of leakage was acceptable as long as the system was running with pure water because purified fresh water was periodically added to the tank. In case of Gd-loaded water, we had to consider environmental issues. There exist no environmental standards for gadolinium anywhere in the world, including in Japan. Consequently, Gd and its compounds are not classified as regulated chemicals by the Water Pollution Prevention Act. However, release of Gd-loaded water to the environment should be avoided because the hazard levels for gadolinium compounds are not yet well established. That is one of the main reasons why we have performed refurbishment work from May 31, 2018, to January 29, 2019. The main tasks are as follows:
@@ 4.4 PMT Exchange and Other Works
During this tank open period, dead PMTs were also exchanged in parallel with sealing work. In total, 136 ID PMTs and 224 OD PMTs were exchanged. For ID PMT replacement, new PMTs under development for Hyper-Kamiokande [HK] were used. In addition to the recovery of SK performance, another purpose of this exchange is to check the performance and long-term stability of HK PMTs in the expected operating environment. Another important task carried out during the tank open work is related to the calibration system inside the tank and studies for more precise detector understanding, such as the replacement of laser injectors and optical fibers, installation of a new light injection system, precise magnetic field mapping, and recording all PMTs dynode directions. All the white Tyvek sheets covering the tank wall were removed for the sealant work, with new Tyvek sheets re-installed afterward.
@@ 5 PURE GD PRODUCTION AND SCREENING
The $\rm Gd_{2}(\rm SO_{4})_{3}\cdot\rm 8H_{2}O$ used for SK-Gd had to meet stringent radiopurity standards to ensure that any added background remained subdominant to those from existing SK solar neutrino and DSNB measurements [3]. Specifically, the total increase in background rates, even after reaching the target concentration of 0.1% Gd in SK, was constrained to be no more than a factor of two.

### v2

@@ 1 INTRODUCTION
While many important measurements have been made with these WC detectors, a major drawback has been their very inefficient response to the presence of thermal neutrons. Now, an affordable and effective technique has been developed and put into practice at Super-Kamiokande to overcome this situation via the addition to water of an optically transparent, low background, low reactivity solute with a large neutron cross section and energetic gamma daughters: gadolinium sulfate octahydrate ($\rm Gd_{2}(\rm SO_{4})_{3}\cdot\rm 8H_{2}O$). See Figure 1.
@@
The idea of loading the famous Super-Kamiokande neutrino and nucleon detector with tons of a water soluble gadolinium compound, a concept which would initially be called GADZOOKS! [Gadolinium Antineutrino Detector Zealously Outperforming Old Kamiokande, Super![1] and then later officially put into practice as SK-Gd, was first introduced at the November 2002 SK Collaboration meeting.
After investigating the use of gadolinium chloride and gadolinium nitrate as our Gd-containing solute, we settled on gadolinium sulfate as the best possible compound for use in Super-K due to its combination of high solubility and high transparency coupled with low reactivity with detector components. The year 2009 saw the funding and construction of EGADS [Evaluating Gadolinium’s Action on Detector Systems], a dedicated, large-scale test facility for gadolinium R&D [2]. This new underground laboratory was built in the Mozumi mine explicitly to demonstrate that the loading of Super-Kamiokande would be safe and effective (see Section 3).
@@ 2.1 Neutron Capture in Water
In pure water, the emitted neutrons thermalize and are captured by protons, with the newly formed deuterons each emitting a single 2.2 MeV gamma ray which can then be measured as delayed signals. Since the fourth phase of SK [SK-IV] started in 2008, new electronics equipment has been introduced [7], making it possible to measure delayed signals. However, in order to achieve the background reduction required for the DSNB analysis mentioned below, the neutron detection efficiency had to be set relatively low (to about 20%). But the situation changes with the introduction of gadolinium. When a neutron is captured by gadolinium, it emits a cascade of gamma rays with a total energy of about 8 MeV, as depicted in Figure 1. This is higher than the energy threshold in SK, and thus the detection efficiency improves drastically even under the same trigger conditions as in the pure-water case. Furthermore, gadolinium has an exceptionally high neutron capture cross section compared to all other elements; it will capture about 75% of thermal neutrons with a modest concentration of only 0.03% Gd by mass dissolved in the water.
@@ 2.2 Diffuse Supernova Neutrino Background
Core-collapse supernovae are powerful explosions that occur in the final stages of the evolution of massive stars, during which roughly 99% of their enormous release of gravitational energy is carried away by neutrinos. Since the beginning of the universe, there have been $\mathcal{O}(10^{18})$ core-collapse supernovae, and the neutrinos released in these explosions fill the current universe as background radiation. This is called the diffuse supernova neutrino background [DSNB]. DSNB observations are expected to provide valuable insights into astrophysics, such as past star formation rates, the frequency of supernova explosions, and the formation of black holes which cannot be seen by optical observations [8, 9, 10, 11]. Since the 1980s, DSNB searches have been conducted using various detectors; however, before the start of SK-Gd no nonzero DSNB flux had been observed. Super-Kamiokande currently has the most stringent constraints on this flux [12, 13, 14], but the expected event rate of DSNB interactions even in the huge SK detector is quite low - at most a few per year - and they are buried under background events, making observation extremely difficult.
@@ 2.3 Galactic Supernova Neutrinos
As mentioned above, many supernovae have occurred to date, but supernova neutrinos have only been observed once. In 1987, neutrinos from a supernova explosion that occurred in the Large Magellanic Cloud, 50 kpc from Earth, were observed at Kamiokande-II, IMB, and Baksan, with 11, 8, and 5 neutrino events, respectively[15, 16, 17]. This observation proved that the basic mechanism of gravitational core collapse that occurs in the final stages of star evolution is correct. Unfortunately, the rate of supernovae close enough to detect neutrinos is estimated at only a few per century [18], so the chance of one occurring within the next decade is low, around 10-30%. However, if such a supernova occurs and a large number of neutrino events are observed, it would be a great advance in our understanding of core-collapse supernovae. Therefore, neutrino observatories around the world are currently preparing for the next nearby supernova to occur.
If a supernova occurs at the center of our galaxy, Super-Kamiokande will detect thousands of neutrinos. Here, the dominant neutrino interaction is the IBD described above, but if a large number of neutrinos are detected, we can also expect detection through other interactions. These include, for example, elastic scattering [ES] : $\nu+e^{-}\rightarrow\nu+e^{-}$; charged current interaction on oxygen : $\nu_{e}+\rm{{}^{16}O}\rightarrow e^{-}+\rm{{}^{16}F}$, $\bar{\nu_{e}}+\rm{{}^{16}O}\rightarrow e^{+}+\rm{{}^{16}N}$; and neutral current interaction on oxygen : $\nu+\rm{{}^{16}O}\rightarrow\nu+\rm{{}^{16}O^{*}}\rightarrow\nu+\rm{{}^{15}O/^{15}N}+\rm{n/p}+\gamma$  [19]. In particular, the direction of electrons produced by ES is strongly correlated with the direction of the original neutrinos, and therefore ES events can be used by Super-Kamiokande to determine the location of the supernova in the sky. Here, it is extremely important to determine the direction to the supernova as quickly as possible. Neutrinos are made before the huge increase in brightness of the dying star, and thus these neutrinos arrive at Earth earlier – anywhere from minutes to hours, depending on the stellar environment and type of progenitor star – than the supernova’s electromagnetic signals. Consequently, if the sky location of an ongoing explosion obtained from neutrino observations can be transmitted quickly enough to the astronomical community, then it should be possible for them to observe the beginning of the shock breakout of a core-collapse supernova.
@@ 2.4 Pre-supernova Neutrinos
As mentioned in the previous section, many neutrinos are expected to be seen by SK-Gd when a core-collapse supernova occurs within our galactic neighborhood, but if the explosion is relatively close to Earth (<<1 kpc) neutrinos could also be observed just before core collapse. Stars during the final stage of their lives are supported by nuclear fusion of heavy nuclei such as silicon, and the cooling mechanism is mainly by neutrino emission via thermal and nuclear processes. These are called pre-supernova neutrinos [21]. Although pre-supernova neutrinos have not yet been found, their observation is very important because this silicon burning phase is thought to last for several days before the core collapse; detecting these neutrinos opens up the possibility of creating a supernova alert. It also would serve to guard against the unfortunate case of SK being down for planned calibration or maintenance just when a nearby star explodes. However, because the flux of these pre-supernova neutrinos is low, the possibility of their measurement is limited to very large, very nearby stars such as Betelgeuse. In fact, Super-Kamiokande has achieved highly efficient neutron capture in SK-Gd and significantly reduced background events, making the detection of neutrinos before a supernova explosion a realistic goal. Since 2021, SK-Gd has been monitoring pre-supernova neutrinos. The sensitivity of the system strongly depends on the supernova model and mass ordering, but in the case of Betelgeuse, for example, it should be possible to issue an alarm between 2 and 12 hours before its core collapse at a significance level corresponding to a false alarm rate of no more than 1 per century [22]. In addition, joint observations with the KamLAND experiment began in May 2023. By integrating the complementary characteristics of the two detectors, the combined alarm system improves its sensitivity to pre-supernova neutrinos. For example, in the most optimistic model, the warning time is 1.5 hours longer than with SK-Gd alone. The combined alert system also reduces the dead time for pre-supernova neutrino detection and facilitates continuous monitoring even if one of the detectors is temporarily offline [23]. [^1] As is currently the case for KamLAND at the time of this writing. That detector is temporarily offline for a major upgrade that started in 2024 and is expected to be completed in 2027.
@@ 2.5.1 Reactor Neutrinos
One of the physics opportunities of SK-Gd is the observation of reactor neutrinos. Since the spectrum of reactor neutrinos is soft, the detection is overwhelmed by radioactive backgrounds and hopeless if we use only the prompt signals of IBD. However, if a delayed neutron signal can be detected, the background events can be drastically reduced, opening the possibility of observing reactor neutrinos. Nevertheless, it is necessary to realize a lower energy threshold for the prompt signal as compared to the DSNB search. This is a long-baseline reactor neutrino observation, famous for measuring neutrino oscillations for $\theta_{12}$ and $\Delta m^{2}_{21}$ in the KamLAND experiment [24]. In SK-Gd, reactor neutrinos are expected to be observed on the order of 100 per year from the currently limited cohort of operating Japanese reactors.
@@ 2.5.2 Atmospheric Neutrinos
In GeV-scale neutrino observations, improving the neutron detection efficiency using Gd is expected to improve the ability to distinguish between neutrinos and antineutrinos. This is because antineutrinos are more likely to produce neutrons in the primary interaction than neutrinos, and the number of neutrons detected after neutrino interactions is expected to reflect this tendency. However, the number of neutrons produced in the final state because of interactions within the nucleus increases with energy, and the neutron multiplicity is much larger than would be predicted simply by considering only primary interactions. Therefore, understanding hadron-nucleus interactions inside the nucleus and in the detector is crucial for this purpose. For atmospheric neutrinos, the ability to distinguish between neutrinos and antineutrinos improves the sensitivity to neutrino mass ordering by taking advantage of the different oscillation patterns of their neutrinos as they pass through the Earth’s core and mantle [25].
@@ 2.5.3 Accelerator Neutrinos
T2K is a long-baseline neutrino experiment in which accelerator neutrinos generated at J-PARC in Tokai, Ibaraki, Japan, are observed at SK, approximately 300 km away. This experiment is expected to also improve its measurement sensitivity by improving the ability to distinguish between neutrinos and antineutrinos, as mentioned in the previous section [26]. Increased neutron detection efficiency is particularly important for antineutrino appearance analysis, as contamination of neutrino events is currently expected to be relatively large. Preliminary studies have shown that this contamination is reduced from 30% to 13% by gadolinium loading. In this study, we have confirmed that, although still very preliminary, the introduction of additional $\bar{\nu}$-enhanced and $\nu$-enhanced samples slightly improves the sensitivity in the analysis of CP violation in the lepton sector.
@@ 3 EGADS DEMONSTRATOR
The EGADS facility includes a 200 ton scale model of SK complete with 240 photomultiplier tubes (227 50-cm SK-style tubes plus an additional 13 prototype tubes that were evaluated for use in the Hyper-Kamiokande project [^2] Hyper-Kamiokande [HK] is the next-generation successor to Super-Kamiokande, a similar but much larger WC detector containing eight times SK’s fiducial volume [27]. HK is currently under construction in a neighboring mountain in Kamioka and is expected to begin data-taking in 2028.), and two different selective water filtration systems. There is also a 15-ton pre-mixing and pre-treatment tank where the $\rm Gd_{2}(\rm SO_{4})_{3}\cdot\rm 8H_{2}O$ is dissolved and cleaned prior to injection into the main 200-ton tank, and a custom-built water transparency measuring system called UDEAL [Underground Device Evaluating Attenuation Length].
@@ 4 TANK REFURBISHMENT FOR GD-LOADING
Since the beginning of Super-K’s operation in 1996, there had been a water leak from the Super-K tank, ranging from 1 to 4 tons per day, depending on the water level in the surrounding rock. That amount of leakage was acceptable as long as the system was running with pure water because purified fresh water was periodically added to the tank. In the case of Gd-loaded water, we had to consider environmental issues. There exist no environmental standards for gadolinium anywhere in the world, including in Japan. Consequently, Gd and its compounds are not classified as regulated chemicals by the Water Pollution Prevention Act. However, release of Gd-loaded water to the environment should be avoided because the hazard levels for gadolinium compounds are not yet well established. That is one of the main reasons why we performed refurbishment work from May 31, 2018, to January 29, 2019. The main tasks were as follows:
@@ 4.4 PMT Exchange and Other Works
During this tank open period, dead PMTs were also exchanged in parallel with sealing work. In total, 136 ID PMTs and 224 OD PMTs were exchanged. For ID PMT replacement, new PMTs under development for Hyper-Kamiokande [HK] were used. In addition to the recovery of SK performance, another purpose of this exchange was to check the performance and long-term stability of HK PMTs in the expected operating environment. Another important task carried out during the tank open work was related to the calibration system inside the tank and studies for more precise detector understanding, such as the replacement of laser injectors and optical fibers, installation of a new light injection system, precise magnetic field mapping, and recording all PMTs’ dynode directions. All the white Tyvek sheets covering the tank wall were removed for the sealant work, with new Tyvek sheets re-installed afterward.
@@ 5 PURE GD PRODUCTION AND SCREENING
The $\rm Gd_{2}(\rm SO_{4})_{3}\cdot\rm 8H_{2}O$ used for SK-Gd had to meet stringent radiopurity standards to ensure that any added background remained subdominant to those from existing SK solar neutrino and DSNB measurements [3]. Specifically, the total increase in background rates, even after reaching the target concentration of 0.1% Gd in SK, was required to be no more than a factor of two.